OpenStudy (anonymous):

Write the equation in standard form of the line through the points (-1,-3) and (2,-1)

OpenStudy (anonymous):

did you find the slope?

OpenStudy (anonymous):

2/3?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes, right 3, up 2 slope is \(\frac{2}{3}\)

OpenStudy (anonymous):

now pick a point and write the point slope formula \[y-y_1=m(x-x_1)\] if you pick \((2,-1)\) you get \[y+1=\frac{2}{3}(x-2)\]

OpenStudy (anonymous):

solve for \(y\) and you aer done

OpenStudy (anonymous):

what is the final answer? @satellite73

OpenStudy (anonymous):

i don't know, i didn't do it

OpenStudy (anonymous):

I am not really sure how to solve it

OpenStudy (anonymous):

\[y+1=\frac{2}{3}(x-2)\]steps are always the same, distribute first, get \[y+1=\frac{2}{3}x-\frac{4}{3}\] then subtract \(1\) from both sides and get \[y=\frac{2}{3}x-\frac{7}{3}\]

OpenStudy (anonymous):

isn't that slope intercept form?

OpenStudy (anonymous):

yes it is what form do you want it in?

OpenStudy (anonymous):

standard

OpenStudy (anonymous):

ooh if i could read "Write the equation in standard form " i would know

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

in that case start with \[y+1=\frac{2}{3}(x-2)\] then multiply both sides by \(3\) to get rid of annoying fraction and get \[3y+3=2(x-2)\] first

OpenStudy (anonymous):

then distribute to get \[3y+3=2x-4\]

OpenStudy (anonymous):

subtract \(3y\) and get \[3=2x-3y+4\] then subtract \(4\) to get \[-1=2x-3y\] or if you prefer \[2x-3y=-1\]